LASER ABLATION FOR SAMPLE PREPARATION IN FOURIER TRANSFORM MICROWAVE SPEC-TROSCOPY

\underline{M. C. L. GERRY}, Department of Chemistry, The University of British Columbia, 6174 University Boulevard, Vancouver, British Columbia, Canada, V6T 1Z3.

After a rather slow start, laser ablation of solids is finding increasingly widespread use in sample preparation for Fourier transform microwave spectroscopy. An account will be given of the use of this technique, with examples of its achievements and suggestions for future applications.